# Smooth Tchebycheff Scalarization for Multi-Objective Optimization

### v2

@@ Abstract
Multi-objective optimization problems can be found in many real-world applications, where the objectives often conflict each other and cannot be optimized by a single solution. In the past few decades, numerous methods have been proposed to find Pareto solutions that represent different optimal trade-offs among the objectives for a given problem. However, these existing methods could have high computational complexity or may not have good theoretical properties for solving a general differentiable multi-objective optimization problem. In this work, by leveraging the smooth optimization technique, we propose a novel and lightweight smooth Tchebycheff scalarization approach for gradient-based multi-objective optimization. It has good theoretical properties for finding all Pareto solutions with valid trade-off preferences, while enjoying significantly lower computational complexity compared to other methods. Experimental results on various real-world application problems fully demonstrate the effectiveness of our proposed method.
@@ 1 Introduction
There are two main types of gradient-based multi-objective optimization algorithms: the scalarization approach (Miettinen, 1999; Ehrgott, 2005) and the adaptive gradient method (Fliege & Svaiter, 2000; Désidéri, 2012). However, both algorithms have their own disadvantages. The simple linear scalarization is the most straightforward method, but will completely miss all the solutions on the non-convex part of the optimal Pareto front (Das & Dennis, 1997; Ehrgott, 2005). The classical Tchebycheff scalarization can find all Pareto solutions that satisfy the decision-maker’s exact trade-off preference (Bowman, 1976; Steuer & Choo, 1983), but it suffers from a slow convergence rate due to its nonsmooth (i.e., non-differentiable) nature. To overcome the limitations of scalarization approaches, many adaptive gradient methods have been proposed in the past few decades (Fliege & Svaiter, 2000; Désidéri, 2012). These methods aim to find a valid gradient direction that improves the performance of all objectives at each iteration. While these methods have theoretical guarantees for finding a Pareto stationary solution, their high pre-iteration computational complexity makes them less suitable for handling large-scale problems in real-world applications.
@@
Instead of proposing another adaptive gradient method, this work revisits the straightforward scalarization approach. In particular, we leverage the powerful smooth approximation technique (Nesterov, 2005; Beck & Teboulle, 2012) to develop a lightweight yet efficient scalarization approach with promising theoretical properties. The main contributions of this work are summarized as follows:  [^1] Our codes will be open-source upon publication.
  • We propose a novel smooth Tchebycheff (STCH) scalarization approach for gradient-based multi-objective optimization, which can serve as a fast alternative to the classic Tchebycheff scalarization.
@@ 2 Preliminaries and Related Work for Multi-Objective Optimization
[…]
@@
[…]
@@
where $\bm{\lambda}=(\lambda_{1},\ldots,\lambda_{m})$ is a preference vector on the simplex $\bm{\Delta}^{m-1}=\{\bm{\lambda}|\sum_{i=1}^{m}\lambda_{i}=1,\lambda_{i}\geq 0\ \forall i\}$ over the $m$ objectives. Once a preference $\bm{\lambda}$ is given, we can obtain a solution by solving the single-objective scalarization problem (4). However, from the perspective of multi-objective optimization, this method can only find solutions on the convex hull of the Pareto front (Boyd & Vandenberghe, 2004; Ehrgott, 2005), and all solutions on the non-convex parts of the Pareto front will be missing (Das & Dennis, 1997). In fact, there is no guarantee that a general scalarization method can find all Pareto solutions, where the Tchebycheff scalarization method is an exception (Miettinen, 1999).
@@
A feasible solution $\bm{x}\in\mathcal{X}$ is weakly Pareto optimal for the original problem 1 if and only if there exists a valid preference vector $\bm{\lambda}$ such that $\bm{x}$ is an optimal solution of the Tchebycheff scalarization problem (5).
@@
According to Désidéri (2012), the obtained $\bm{d}_{t}$ should either be a valid gradient direction to improve all objectives or $\bm{d}_{t}=\bm{0}$ which means the solution $\bm{x}_{t}$ is Pareto stationary. Therefore, we can use a simple gradient-based method to find a Pareto stationary solution (Fliege et al., 2019). In recent years, many work have adopted these gradient-based algorithms to tackle multi-objective optimization problems in the machine learning community (Sener & Koltun, 2018; Lin et al., 2019; Mahapatra & Rajan, 2020; Ma et al., 2020; Momma et al., 2022). Similar adaptive gradient methods are also developed to solve multi-task learning problems (Yu et al., 2020; Liu et al., 2021a, b, 2022; Navon et al., 2022; Senushkin et al., 2023; Lin et al., 2023). Stochastic multi-objective optimization is another important topic for real-world applications (Liu & Vicente, 2021; Zhou et al., 2022; Fernando et al., 2023; Chen et al., 2023; Xiao et al., 2023).
However, at each iteration, these adaptive gradient algorithms typically need to calculate the gradients for each objective separately (e.g., via total $m$ different backpropagation) and solve a quadratic programming problem such as (7). This high pre-iteration complexity will lead to high computational overhead for solving large-scale problems such as training a deep neural network model.
@@ 3.2 Smooth Tchebycheff Scalarization
[…]
@@
[…]
@@
[…]
@@
[…]
@@ 3.3 Properties for Multi-Objective Optimization
We can prove this theorem by analyzing the geometry relation between the level surface of STCH scalarization and the Pareto front. The key step is to find a theoretical upper bound of $\mu^{*}$ such that the level surface of STCH scalarization with any $\mu<\mu^{*}$ is totally enclosed by the Pareto front. Details can be found in Appendix A.2. ∎
@@
[…]
@@ 3.4 STCH Scalarization for Pareto Set Learning
Classic scalarization and adaptive gradient methods all focus on finding a single or finite set of Pareto solutions, but the whole Pareto set could be an $(m-1)$-dimensional manifold that contains infinite solutions (Hillermeier, 2001). Recently, some works have been proposed to build a set model to approximate the whole Pareto set or front (Parisi et al., 2014; Dosovitskiy & Djolonga, 2019; Lin et al., 2020; Navon et al., 2021; Ruchte & Grabocka, 2021; Lin et al., 2022b, c; Chen & Kwok, 2022; Jain et al., 2023; Zhang et al., 2023). A typical Pareto set model can be defined as:
@@
Meannormal-$\downarrow$ MEDnormal-$\downarrow$ 11.25normal-$\uparrow$ 22.5normal-$\uparrow$ 30normal-$\uparrow$
@@ 4 Experimental Studies
In this section, we evaluate our proposed STCH scalarization approach on a diverse set of multi-objective optimization problems, which includes finding a single or a few Pareto solutions, as well as learning the whole Pareto set.
@@ 4.1 Multi-Task Learning
We compare the STCH scalarization with the single-task learning (STL) baseline and (1) other scalarization methods: equal weight (EW), geometric mean loss (GLS) (Chennupati et al., 2019), random weights (RLW) (Lin et al., 2022a), and classical Tchebycheff scalarization (TCH); (2) adaptive loss methods: UW (Kendall et al., 2018), DWA (Liu et al., 2019), IMTL-L, and IGBv2 (Dai et al., 2023); (3) adaptive gradient methods: MGDA (Sener & Koltun, 2018), GradNorm (Chen et al., 2018), PCGrad (Yu et al., 2020), GradDrop (Chen et al., 2020), GradVac (Wang & Tsvetkov, 2021), IMTL-G, CAGrad (Liu et al., 2021a), MTAdam (Malkiel & Wolf, 2021), Nash-MTL (Navon et al., 2022), MetaBalance (He et al., 2022), MoCo (Fernando et al., 2023), Alighed-MTL (Senushkin et al., 2023), and DB-MTL (Lin et al., 2023); (4) hybrid loss-gradient balancing method: IMTL (Liu et al., 2021b). All methods are implemented with the LibMTL library (Lin & Zhang, 2023). The problem and model details can be found in Appendix C.1.
@@
Office-31 (Saenko et al., 2010) is an image classification dataset across $3$ domains (Amazon, DSLR, and Webcam). According to Table 4, STCH scalarization can achieve the best overall performance among all methods. In fact, it is the only method that can dominate the STL baseline on all tasks. The full results can be found in Appendix D.1.
QM9 (Ramakrishnan et al., 2014) is a molecular property prediction dataset with 11 tasks. The results in Table 4 suggest that STCH scalarization has a very-close-to-best overall $\Delta_{p}$ with DB-MTL, while enjoying a much faster run time. The full results can be found in Appendix D.2.
@@
4.2 Efficient Pareto Set Learning
Baseline Methods
Evaluation Metric
Following the related work, we use the hypervolume difference ($\Delta\text{HV}$(Zitzler et al., 2007) to measure the qualities of approximate Pareto fronts learned by different methods. A lower $\Delta\text{HV}$ means that the learned Pareto front is closer to the ground truth Pareto front and hence has better performance. A detailed definition of $\Delta\text{HV}$ can be found in Appendix C.2.3.
Results and Analysis
We evaluate all methods on six synthetic benchmark problems (F1-F6) as well as five real-world engineering design problems for Bar Truss, Hatch Cover, Disk Brake, Gear Train, and Rocket Injector. According to Table 3, the Pareto set learning model with STCH scalarization performs the best on $10$ out of $11$ problems. The approximate Pareto fronts learned by different scalarization methods for the rocket injector design are shown in Figure 4. The Pareto front learned by STCH can find better trade-offs than others. These results demonstrate that STCH is a promising scalarization approach for Pareto set learning. Problem details can be found in Appendix C.2.1 and C.2.2.
5 Conclusion and Discussion
Conclusion
In this work, we have proposed a novel smooth Tchebycheff (STCH) scalarization approach for efficient gradient-based multi-objective optimization. It has a fast convergence rate and low pre-iteration complexity, while enjoying many good theoretical properties. Experimental studies also show that it can obtain very promising performance on various multi-objective optimization problems. We believe that the proposed STCH scalarization can serve as a powerful yet straightforward method to tackle differentiable multi-objective optimization.

### v3

@@ Abstract
Multi-objective optimization problems can be found in many real-world applications, where the objectives often conflict each other and cannot be optimized by a single solution. In the past few decades, numerous methods have been proposed to find Pareto solutions that represent optimal trade-offs among the objectives for a given problem. However, these existing methods could have high computational complexity or may not have good theoretical properties for solving a general differentiable multi-objective optimization problem. In this work, by leveraging the smooth optimization technique, we propose a lightweight and efficient smooth Tchebycheff scalarization approach for gradient-based multi-objective optimization. It has good theoretical properties for finding all Pareto solutions with valid trade-off preferences, while enjoying significantly lower computational complexity compared to other methods. Experimental results on various real-world application problems fully demonstrate the effectiveness of our proposed method.
@@ 1 Introduction
There are two main types of gradient-based multi-objective optimization algorithms: the scalarization approach (Miettinen, 1999; Ehrgott, 2005) and the adaptive gradient method (Fliege et al., 2019). However, both algorithms have their own disadvantages. The simple linear scalarization is the most straightforward method, but will completely miss all the solutions on the non-convex part of the optimal Pareto front (Das & Dennis, 1997; Ehrgott, 2005). The classical Tchebycheff scalarization can find all Pareto solutions that satisfy the decision-maker’s exact trade-off preference (Bowman, 1976; Steuer & Choo, 1983), but it suffers from a slow convergence rate due to its nonsmooth (i.e., non-differentiable) nature. To overcome the limitations of scalarization approaches, many adaptive gradient methods have been proposed in the past few decades (Fliege & Svaiter, 2000; Schäffler et al., 2002; Désidéri, 2012). These methods aim to find a valid gradient direction that improves the performance of all objectives at each iteration. While these methods have theoretical guarantees for finding a Pareto stationary solution, their high pre-iteration computational complexity makes them less suitable for handling large-scale problems in real-world applications.
@@
Instead of proposing another adaptive gradient method, this work revisits the straightforward scalarization approach. In particular, we leverage the powerful smooth approximation technique (Nesterov, 2005; Beck & Teboulle, 2012) to develop a lightweight yet efficient scalarization approach with promising theoretical properties. The main contributions of this work are summarized as follows.  [^1] Our source code is available at: github.com/Xi-L/STCH.
  • We propose a smooth Tchebycheff (STCH) scalarization approach for gradient-based multi-objective optimization, which can serve as a fast alternative to the classic Tchebycheff scalarization.
@@ 2 Preliminaries and Related Work for Multi-Objective Optimization
Let $\bm{x}^{(a)},\bm{x}^{(b)}\in\mathcal{X}$ be two solutions for problem (1), $\bm{x}^{(a)}$ is said to dominate $\bm{x}^{(b)}$, denoted as $\bm{f}(\bm{x}^{(a)})\prec\bm{f}(\bm{x}^{(b)})$, if and only if $f_{i}(\bm{x}^{(a)})\leq f_{i}(\bm{x}^{(b)})\ \forall i\in\{1,...,m\}$ and $f_{j}(\bm{x}^{(a)})<f_{j}(\bm{x}^{(b)})\ \exists j\in\{1,...,m\}$. In addition, $\bm{x}^{(a)}$ is said to strictly dominate $\bm{x}^{(b)}$ (i.e., $\bm{f}(\bm{x}^{(a)})\prec_{\text{strict}}\bm{f}(\bm{x}^{(b)})$), if and only if $f_{i}(\bm{x}^{(a)})<f_{i}(\bm{x}^{(b)})\ \forall i\in\{1,...,m\}$.
@@
A solution $\bm{x}^{\ast}\in\mathcal{X}$ is Pareto optimal if there is no $\bm{x}\in\mathcal{X}$ such that $\bm{f}(\bm{x})\prec\bm{f}(\bm{x}^{\ast})$. A solution $\bm{x}^{\prime}\in\mathcal{X}$ is weakly Pareto optimal if there is no $\bm{x}\in\mathcal{X}$ such that $\bm{f}(\bm{x})\prec_{\text{strict}}\bm{f}(\bm{x}^{\prime})$.
@@
where $\bm{\lambda}=(\lambda_{1},\ldots,\lambda_{m})$ is a preference vector on the simplex $\bm{\Delta}^{m-1}=\{\bm{\lambda}|\sum_{i=1}^{m}\lambda_{i}=1\text{ and }\lambda_{i}\geq 0,\forall i\}$ over the $m$ objectives. Once a preference $\bm{\lambda}$ is given, we can obtain a solution by solving the single-objective scalarization problem (4). However, from the perspective of multi-objective optimization, this method can only find solutions on the convex hull of the Pareto front (Boyd & Vandenberghe, 2004; Ehrgott, 2005), and all solutions on the non-convex parts of the Pareto front will be missing (Das & Dennis, 1997). In fact, there is no guarantee that a general scalarization method can find all Pareto solutions, where the Tchebycheff scalarization method is an exception (Miettinen, 1999).
@@
A feasible solution $\bm{x}\in\mathcal{X}$ is weakly Pareto optimal for the original problem (1) if and only if there exists a valid preference vector $\bm{\lambda}$ such that $\bm{x}$ is an optimal solution of the Tchebycheff scalarization problem (5).
@@
According to Désidéri (2012), the obtained $\bm{d}_{t}$ should either be a valid gradient direction to improve all objectives or $\bm{d}_{t}=\bm{0}$ which means the solution $\bm{x}_{t}$ is Pareto stationary. Therefore, we can use a simple gradient-based method to find a Pareto stationary solution (Fliege et al., 2019). In recent years, many work have adopted these gradient-based algorithms to tackle multi-objective optimization problems in the machine learning community (Sener & Koltun, 2018; Lin et al., 2019; Mahapatra & Rajan, 2020; Ma et al., 2020; Momma et al., 2022). Similar adaptive gradient methods are also developed to solve multi-task learning problems (Yu et al., 2020; Liu et al., 2021a, b, 2022; Navon et al., 2022; Senushkin et al., 2023; Lin et al., 2023a). Stochastic multi-objective optimization is another important topic for real-world applications (Liu & Vicente, 2021; Zhou et al., 2022; Fernando et al., 2023; Chen et al., 2023; Xiao et al., 2023).
However, at each iteration, these adaptive gradient algorithms typically need to calculate the gradients for each objective separately (e.g., via total $m$ different backpropagation) and solve a quadratic programming problem such as (7). This high pre-iteration complexity will lead to high computational overhead for solving large-scale problems such as training a deep neural network model. In addition, some recent works show that many adaptive gradient methods cannot significantly outperform a well-tuned or even random linear scalarization for deep multi-task learning problems (Kurin et al., 2022; Xin et al., 2022; Lin et al., 2022a; Royer et al., 2023), although linear scalarization fails to fully explore the whole Pareto front (Hu et al., 2023).
@@ 3.2 Smooth Tchebycheff Scalarization
The smooth Tchebycheff scalarization $g^{(\text{STCH})}_{\mu}(\bm{x}|\bm{\lambda})$ is a smoothing function of the classical Tchebycheff scalarization $g^{(\text{TCH})}(\bm{x}|\bm{\lambda})$ that satisfies Definition 3.2.
@@
$g^{(\text{STCH})}_{\mu}(\bm{x}|\bm{\lambda})-\mu\log{m}\leq g^{(\text{TCH})}(\bm{x}|\bm{\lambda})\leq g^{(\text{STCH})}_{\mu}(\bm{x}|\bm{\lambda}).$ (10)
@@
If all objective functions $\{f_{1}(\bm{x}),\ldots,f_{m}(\bm{x})\}$ are convex, then the STCH scalarization $g^{(\text{STCH})}_{\mu}(\bm{x}|\bm{\lambda})$ with any valid $\mu$ and $\bm{\lambda}$ is convex.
@@
If all objective functions are convex, with a properly chosen smoothing parameter $\mu$, we can obtain an $\epsilon$-optimal solution to the nonsmooth $g^{(\text{TCH})}_{\mu}(\bm{x}|\bm{\lambda})$ within $O(\frac{1}{\epsilon})$ iterations by solving the smooth counterpart $g^{(\text{STCH})}_{\mu}(\bm{x}|\bm{\lambda})$.
@@ 3.3 Properties for Multi-Objective Optimization
We can prove this theorem by analyzing the geometry relation between the level surface of STCH scalarization and the Pareto front. The key step is to find a theoretical upper bound of $\mu^{*}$ such that the level surface of STCH scalarization with any $\mu<\mu^{*}$ is totally enclosed by the Pareto front. The mild conditions are formally defined in Assumption A.1, and detailed proof can be found in Appendix A.2. ∎
@@
If there exists a solution ^$\hat{}\bm{x}$ such that $\nabla g^{(\text{STCH})}_{\mu}(\hat{}\bm{x}|\bm{\lambda})=0$, then ^$\hat{}\bm{x}$ is a Pareto stationary solution of the original multi-objective optimization problem (1).
@@ 3.4 STCH Scalarization for Pareto Set Learning
Classic scalarization and adaptive gradient methods all focus on finding a single or finite set of Pareto solutions, but the whole Pareto set could be an $(m-1)$-dimensional manifold that contains infinite solutions (Hillermeier, 2001). Recently, some works have been proposed to build a model to approximate the Pareto set or front (Parisi et al., 2014; Dosovitskiy & Djolonga, 2019; Lin et al., 2020b; Navon et al., 2021; Ruchte & Grabocka, 2021; Lin et al., 2022b, c; Chen & Kwok, 2022; Jain et al., 2023; Lin et al., 2023b; Zhang et al., 2023). A Pareto set model can be defined as:
@@
Single Task Baseline
[…]
Adaptive Gradient Method
@@ 4 Experimental Studies
We evaluate the proposed STCH scalarization approach on finding a single Pareto solution and the whole Pareto set.
@@ 4.1 Multi-Task Learning
We compare the STCH scalarization with the single-task learning (STL) baseline and (1) other scalarization methods: equal weight (EW), geometric mean loss (GLS) (Chennupati et al., 2019), random weights (RLW) (Lin et al., 2022a), and classical Tchebycheff scalarization (TCH); (2) adaptive loss methods: UW (Kendall et al., 2018), DWA (Liu et al., 2019), IMTL-L, and IGBv2 (Dai et al., 2023); (3) adaptive gradient methods: MGDA (Sener & Koltun, 2018), GradNorm (Chen et al., 2018), PCGrad (Yu et al., 2020), GradDrop (Chen et al., 2020), GradVac (Wang & Tsvetkov, 2021), IMTL-G, CAGrad (Liu et al., 2021a), MTAdam (Malkiel & Wolf, 2021), Nash-MTL (Navon et al., 2022), MetaBalance (He et al., 2022), MoCo (Fernando et al., 2023), Alighed-MTL (Senushkin et al., 2023), and DB-MTL (Lin et al., 2023a); (4) hybrid loss-gradient balancing method: IMTL (Liu et al., 2021b). All methods are implemented with the LibMTL library (Lin & Zhang, 2023). The problem and model training details can be found in Appendix D.1.
@@
Office-31 (Saenko et al., 2010) is an image classification dataset across $3$ domains (Amazon, DSLR, and Webcam). According to Table 4, STCH scalarization can achieve the best overall performance among all methods. In fact, it is the only method that can dominate the STL baseline on all tasks. The full results can be found in Appendix E.1.
QM9 (Ramakrishnan et al., 2014) is a molecular property prediction dataset with 11 tasks. The results in Table 4 suggest that STCH scalarization has a very-close-to-best overall $\Delta_{p}$ with DB-MTL, while enjoying a much faster run time. The full results can be found in Appendix E.2.
@@
4.2 Efficient Pareto Set Learning
Baseline Methods
Evaluation Metric
Following the related work, we use the hypervolume difference ($\Delta\text{HV}$(Zitzler et al., 2007) to measure the qualities of approximate Pareto fronts learned by different methods. A lower $\Delta\text{HV}$ means that the learned Pareto front is closer to the ground truth Pareto front and hence has better performance. A detailed definition of $\Delta\text{HV}$ can be found in Appendix D.2.3.
Results and Analysis
We evaluate all methods on six synthetic benchmark problems (F1-F6) as well as five real-world engineering design problems for Bar Truss, Hatch Cover, Disk Brake, Gear Train, and Rocket Injector. According to Table 3, the Pareto set learning model with STCH scalarization performs the best on $10$ out of $11$ problems. The approximate Pareto fronts learned by different scalarization methods for the rocket injector design are shown in Figure 4. The Pareto front learned by STCH can find better trade-offs than others. These results demonstrate that STCH is a promising scalarization approach for Pareto set learning. Problem details can be found in Appendix D.2.1 and D.2.2.
Table 5: The results of STCH with different smooth parameters $\mu$ on $5$ real-world engineering desgin problems.
BarTruss HatchCover DiskBrake GearTrain RocketInjector
TCH 9.05e-03 1.01e-02 3.78e-02 3.91e-03 2.73e-02
STCH($\mu=0.01$) 6.94e-03 8.47e-03 3.12e-02 3.79e-03 1.95e-02
STCH($\mu=0.1$) 6.85e-03 8.76e-03 2.79e-02 3.17e-03 2.00e-02
STCH($\mu=0.5$) 6.21e-03 7.81e-03 2.96e-02 3.36e-03 2.24e-02
STCH($\mu=1$) 6.95e-03 8.02e-03 3.32e-02 3.62e-03 3.24e-02
4.3 Effect of Different Smooth Levels
The parameter $\mu$ controls the smoothing level of the STCH scalarization, which acts as a hyperparameter in our proposed method. In this work, we do not aggressively tune $\mu$ for each problem. To investigate the effect of $\mu$, we conduct a new experiment to test STCH with different smooth levels $\mu$ for the multi-objective engineering design problem. As shown in Table 5, STCH scalarization with a small $\mu$ works reasonably well, but there is no single $\mu$ that can achieve the best performance for all problems.
The smooth parameter $\mu$ does have an impact on the convergence rate of the gradient-based algorithm. Roughly speaking, a larger $\mu$ can yield a lower iteration complexity for the smoothed function, while a small enough $\mu$ is required to obtain a good approximate solution for the original non-smooth problem. To better leverage this trade-off, a few homotopy methods (Allen-Zhu & Hazan, 2016; Xu et al., 2016) have been proposed to gradually decrease $\mu$ in a stage-wise manner for the Nesterov smoothing approach (Nesterov, 2005). This method can reduce the iteration complexity from $O(1/\epsilon)$ to $\tilde{O}(1/\epsilon^{1-\theta})$ where $\theta\in(0,1]$ represents the local sharpness around the optimal solution.
However, the homotopy schedule could be problem-specific and should be tuned for each new problem. We do not observe a clear performance improvement by using the homotopy method and, therefore, simply use a fixed $\mu$ in all our experiments. How to adaptively adjust the smooth parameter $\mu$ for each problem during the optimization process could be an interesting research direction in future work.
5 Local and Global Optimality Guarantee
Although Theorem 3.8 provides a promising sufficient condition for STCH scalarization to find all Pareto solutions, the simple gradient-based algorithm can only guarantee to find Pareto stationary solutions for a general multi-objective optimization problem as in Theorem 3.10. We would like to make the following remarks on this theoretical guarantee:
Similar Guarantee with the Other Methods
The global optimality guarantee for STCH is on par with the optimality guarantee for the classic TCH scalarization counterpart as in Theorem 2.3. Every Pareto optimal solution can be found by TCH scalarization, but the subgradient method can only guarantee to find a Pareto stationary solution. This guarantee is also on par with the adaptive gradient methods that are popular in the current literature, but our proposed method has a much lower pre-iteration complexity. Under a stronger assumption, such as all objectives are convex (and hence the STCH function is convex), the proposed method can find the global optimal solution.
Advanced Method for Global Optimality
It could be extremely hard, if not impossible, to provide a convergence guarantee to find the global optimal solution for a general non-convex optimization problem. Some advanced global optimization methods, such as homotopy/graduated optimization (Dunlavy & O’Leary, 2005; Hazan et al., 2016), could be helpful, but it is still a very difficult open problem for gradient-based optimization. We will explore this direction in future work. On the other hand, a local optimal solution is usually good enough for deep learning applications. According to the experimental results, our proposed method achieves promising performance on the MTL problems, while enjoying a significantly faster runtime than those adaptive gradient methods.
6 Conclusion, Limitation, and Future Work
Conclusion
In this work, we have proposed a novel smooth Tchebycheff (STCH) scalarization approach for efficient gradient-based multi-objective optimization. It has a fast convergence rate and low pre-iteration complexity, while enjoying many good theoretical properties. Experimental studies also show that STCH scalarization can obtain promising performance on various multi-objective optimization problems. We believe it can serve as a powerful yet straightforward method to tackle differentiable multi-objective optimization.